4524885555444155-2846657044580587.9
1
Expert's answer
2020-12-15T18:41:48-0500

Let's represent the number "4524885555444155" as the sum of several terms.

"4524885555444155 =4*10^{15}+5*10^{14}+2*10^{13}+"

"+4*10^{12} + 8*10^{11}+8*10^{10}+5*10^{9}+5*10^{8}+"

"+5*10^7+5*10^6+4*10^5+4*10^4+4*10^3+"

"+1*10^2 + 5*10^1 + 5*10^0 + 0*10^{-1}."

Let's represent the number "2846657044580587.9" as the sum of several terms.

"2846657044580587.9= 2*10^{15}+8*10^{14}+4*10^{13}+"

"+6*10^{12}+6*10^{11}+5*10^{10}+7*10^{9}+0*10^8+"

"+4*10^7+4*10^6+5*10^5+8*10^4+0*10^3+"

"+5*10^2+8*10^1+7*10^0+9*10^{-1}."

Now we can rewrite the difference:

"4524885555444155-2846657044580587.9="

"10^{15}*(4-2)+10^{14}*(5-8) + 10^{13}*(2-4)+"

"+10^{12}*(4-6)+10^{11}*(8-6)+10^{10}*(8-5)+"

"+10^9*(5-7)+10^8*(5-0)+10^7*(5-4)+"

"+10^6*(5-4)+10^5*(4-5)+10^4*(4-8)+"

"+10^3*(4-0)+10^2*(1-5)+10^1*(5-8)+"

"+10^0*(5-7)+10^{-1}*(0-9)="

"10^{15}*2 +10^{14}*(-3)+10^{13}*(-2)+10^{12}*(-2)+"

"+10^{11}*2+10^{10}*3+10^9*(-2)+10^8*5+10^7*1+"

"+10^6*1+10^5*(-1)+10^4*(-4)+10^3*4+10^2*(-4)+"

"+10^1*(-3)+10^0*(-2)+10^{-1}*(-9)="

"10^{15}*1 + 10^{14}*(10-3)+10^{13}*(-2)+10^{12}*(-2)+"

"+10^{11}*2+10^{10}*2+10^9*(10-2)+10^8*5+10^7*1+"

"+10^6*0+10^5*(10-1)+10^4*(-4)+10^3*3+10^2*(10-4)+"

"+10^1*(-3)+10^0*(-2)+10^{-1}*(-9)="

"10^{15}*1 + 10^{14}*7+10^{13}*(-2)+10^{12}*(-2)+"

"+10^{11}*2+10^{10}*2+10^9*8+10^8*5+10^7*1+"

"+10^6*0+10^5*9+10^4*(-4)+10^3*3+10^2*6+"

"+10^1*(-3)+10^0*(-2)+10^{-1}*(-9)="

"10^{15}*1 + 10^{14}*6+10^{13}*(10-2)+10^{12}*(-2)+"

"+10^{11}*2+10^{10}*2+10^9*8+10^8*5+10^7*1+"

"+10^6*0+10^5*8+10^4*(10-4)+10^3*3+10^2*5+"

"+10^1*(10-3)+10^0*(-2)+10^{-1}*(-9)="

"10^{15}*1 + 10^{14}*6+10^{13}*8+10^{12}*(-2)+"

"+10^{11}*2+10^{10}*2+10^9*8+10^8*5+10^7*1+"

"+10^6*0+10^5*8+10^4*6+10^3*3+10^2*5+"

"+10^1*7+10^0*(-2)+10^{-1}*(-9)="

"10^{15}*1 + 10^{14}*6+10^{13}*7+10^{12}*(10-2)+"

"+10^{11}*2+10^{10}*2+10^9*8+10^8*5+10^7*1+"

"+10^6*0+10^5*8+10^4*6+10^3*3+10^2*5+"

"+10^1*6+10^0*(10-2)+10^{-1}*(-9)="

"10^{15}*1 + 10^{14}*6+10^{13}*7+10^{12}*8+"

"+10^{11}*2+10^{10}*2+10^9*8+10^8*5+10^7*1+"

"+10^6*0+10^5*8+10^4*6+10^3*3+10^2*5+"

"+10^1*6+10^0*8+10^{-1}*(-9)="

"10^{15}*1 + 10^{14}*6+10^{13}*7+10^{12}*8+"

"+10^{11}*2+10^{10}*2+10^9*8+10^8*5+10^7*1+"

"+10^6*0+10^5*8+10^4*6+10^3*3+10^2*5+"

"+10^1*6+10^0*7+10^{-1}*(10-9)="

"10^{15}*1 + 10^{14}*6+10^{13}*7+10^{12}*8+"

"+10^{11}*2+10^{10}*2+10^9*8+10^8*5+10^7*1+"

"+10^6*0+10^5*8+10^4*6+10^3*3+10^2*5+"

"+10^1*6+10^0*7+10^{-1}*1="

"1 678 228 510 863 567.1"